# Ambitions for theory in the physics of life

### v1

@@ 4.2 Ingredients of a genetic network
$h>0$ means that the TF activates expression and $K$ is the concentration at which this effects is half–maximal. More quantitatively $h$ is a measure of sensitivity. We can rationalize this form by imagining the $h$ molecules of the TF bind cooperatively to sites along the DNA, and this binding influences transcription. To complete our description, and the calculation of the information transmission, we need to know the noise level $\sigma_{g}(c)$.
@@
Figure 23: Schematic of concentration measurements [56, 57]. A receptor of linear dimension $a$ samples a volume $\sim a^{3}$ and thus counts a mean number of molecules $\bar{n}\sim ca^{3}$, where $c$ is the concentration. Molecules random walk in and out of the sensitive volume with a diffusion constant $D$, corresponding to an exchange or correlation time $\tau_{c}\sim a^{2}/D$.
@@
If there is just one receptor of linear dimension $a_{\rm r}$ then the Berg–Purcell argument should still work. If there are $N_{\rm r}$ of these receptors then it is plausible that noise is reduced by a factor $1/\sqrt{N_{\rm r}}$, which is equivalent to saying they act together as a single receptor with effective size $a_{\rm eff}\sim N_{\rm r}a_{\rm r}$. But if the number of receptor becomes large enough that the area occupied by the receptors becomes comparable to the area of the cell surface, $N_{\rm r}a_{\rm r}^{2}\sim a_{\rm cell}^{2}$, then it seems like the whole cell should act as one big receptor and $a_{\rm eff}\rightarrow a_{\rm cell}$. How does this work?
In a fabulous bit of hand waving plus analogies, Berg and Purcell argued that the tendency of random walk trajectories to bounce along a surface leads to correlations in the encounters of individual molecules with nearby receptors. With $N_{\rm r}$ receptors scattered over the surface of the cell, this lead to an effective size
@@
What the Berg–Purcell limit tells us is that the random arrival of molecules at their target site generate an effective concentration noise with variance
@@
Figure 24: Optimizing information about the concentration of a single transcription factor [200]. (A) A map of $F$ from Eq (138) as a function of $K$ and $h$, showing a single clear optimum at reasonable parameters ($c_{\rm max}/c_{0}=1$). (B–F) Optimal input/output relations $\bar{g}_{\rm i}(c)$ with five target genes, with increasing values of $c_{\rm max}/c_{0}$ indicated by dashed vertical lines. At smaller $c_{\rm max}/c_{0}$ the optimal solutions are redundant, so fewer distinct input/output relations are visible. (G) Optimal values of $K/c_{0}$, showing a sequence of bifurcations as $c_{\rm max}/c_{0}$ increases and responses become distinct.
@@
Figure 25: Redundancy is reduced in a system where a single transcription factor (concentration $c$) activates multiple genes $\{g_{1},g_{2},g_{3}\}$ that repress one another, as shown at left. At right are the mean expression levels that optimize information transmission, with $c_{\rm max}/c_{0}=10$. Redrawn from Ref [201].
@@
Figure 26: Schematic of direct transcriptional regulation (DTR) and indirect translational regulation (ITR) [205]. (A) In DTR, as above, activator TFs (green squares at concentration $c$) interact with binding sites along the DNA to activate expression of the regulated gene $g$. (B) In ITR scenario, input molecules (green squares) bind to mRNAs $m$ of protein $y$ (red chain) to make the mRNA inaccessible for translation (gray oval). Translation proceeds from unbound mRNA molecules, giving rise to proteins $y$ (red stars). These proteins act as repressors for gene $g$; the overall mapping from $c\rightarrow g$ is thus activating in both scenarios.
In Figures 24B–F we see that even as the different target genes acquire distinct input/output relations, the optimal solutions still do not make much use of the lowest accessible concentrations; this regime is simply too noisy to allow reliable information transmission. But suppose that the cell makes many mRNA molecules of some intermediate protein $y$, and that the protein $c$ binds to these mRNA and represses translation, as in Fig 26; $y$ can then act indirectly as a transcriptional repressor so that the whole path $c\rightarrow g$ is activating. Then the “measurement” of low concentrations is made in parallel at many sites, rather than at just one site along the DNA, and the response is largest and most reliable at the lowest concentrations. This suggests that a molecule which functions both as a transcriptional activator and a translational repressor could make better use of its full dynamic range. It requires a careful calculation to show that information transmission really can be larger even when we constrain the total number of molecules, but it does [205]. One of the primary maternal morphogens in the fly embryo belongs to a whole class of proteins that function as such dual transcription/translation regulators [206, 207, 208, 209], and our best estimates put this system in the regime where dual regulation enhances information transmission. I found it remarkable that this almost baroque level of biological complexity can be derived as part of the solution to a fundamental physics problem faced by the cell.
Each of these pieces—repression to reduce redundancy, feedback to average over time, diffusion to average over space, and dual regulations—is an interesting problem by itself, and it was possible to make progress with a combination of analytic and (modest) numerical approaches. But all of these things are happening at once in the fly embryo, even just in the gap gene network. To put (most of) these pieces together requires a more sophisticated approach [210].
@@ 4.3 The gap genes, once more
One attractive feature of the MWC model is that it generalizes to having multiple TFs converge to regulate a single gene. Consider that gene $\rm i$ has $H_{\rm ij}$ binding sites for the protein encoded by gene $\rm j$ and $H_{{\rm i}\alpha}$ sites for the maternal input $\alpha$, Then if we follow the same equilibrium statistical mechanics arguments as above, we will find
@@
Figure 29: Maximizing positional information in the gap gene network [210]. Random initial parameter settings produce zero positional information, with all gap genes fully on or off across the length of the embryo. Across $\sim 20,000$ steps of optimization we find parameter settings that allow first one then two, three, and finally all four gap genes to be modulated by the dynamic range of maternal inputs ($1\rightarrow 2\rightarrow 3\rightarrow 4$). The majority of the optimization run then tunes the parameters to maximize the positional information, arriving at a setting ($5$) that behaves very similarly to the real network.
@@
  • If we eliminate feedback loops we can still reach similar levels of positional information but only with unphysically large value of the $H$ parameters; within realistic settings feedback is essential.
@@
I suspect we are just scratching the surface. Perhaps most intriguing observation is that different local optima have broadly consistent gap gene expression patterns, but with shifts and swaps similar to those seen in related species [220, 221, 222].
@@ 5 Conclusion
As physicists we are used to the idea that macroscopic phenomena are described by coarse–grained models, and that in these models many microscopic details have disappeared (see Approach 3 in §1.2). But in the examples we understand, from the inanimate world, we can do this coarse–graining explicitly to see which (relevant) elements of the model survive to influence macroscopic dynamics and which (irrelevant) details are erased. As far as I know nobody knows how to start with a realistic description of ion channels and coarse–grain to arrive at a model of neural networks. Further, we do know that some molecular details must matter, because particular classes of cells in the brain will switch from using one kind of channel to another at crucial moments in development, and of course different types of cells use different combinations of channels. We also know that not all macroscopic quantities are universal, and there is the danger that what is universal might not be relevant for the organism.
@@ Acknowledgements
In addition to the handful of friends who were at the origins of these ideas, the specific things I discussed in these lectures involved collaborations with Naama Brenner, Curtis Callan, Julien Dubuis, Adrienne Fairhall, Dmitry Krotov, Geoff Lewen, Geoff Owen, Mariela Petkova, Marc Potters, Fred Rieke, Shiva Sinha, Thomas Sokolowski, and Aleksandra Walczak. This bland listing doesn’t do justice to all the fun we have had together.

### v2

@@ 4.2 Ingredients of a genetic network
$h>0$ means that the TF activates expression and $K$ is the concentration at which this effect is half–maximal. More quantitatively $h$ is a measure of sensitivity. We can rationalize this form by imagining the $h$ molecules of the TF bind cooperatively to sites along the DNA, and this binding influences transcription. To complete our description, and the calculation of the information transmission, we need to know the noise level $\sigma_{g}(c)$.
@@
Figure 23: Understanding the physical limits to concentration measurements [56, 57]. A receptor of linear size $a$ samples a volume $\sim a^{3}$ and thus counts a mean number of molecules $\bar{n}\sim ca^{3}$, where $c$ is the concentration. Molecules move randomly in and out of the sensitive volume with a diffusion constant $D$, setting the correlation time $\tau_{c}\sim a^{2}/D$. Figure adapted from Ref [57], with permission.
@@
If there is just one receptor of linear dimension $a_{\rm r}$ then the Berg–Purcell argument should still work. If there are $N_{\rm r}$ of these receptors then it is plausible that noise is reduced by a factor $1/\sqrt{N_{\rm r}}$, which is equivalent to saying they act together as a single receptor with effective size $a_{\rm eff}\sim N_{\rm r}a_{\rm r}$. But if the number of receptors becomes large enough that the area occupied by the receptors becomes comparable to the area of the cell surface, $N_{\rm r}a_{\rm r}^{2}\sim a_{\rm cell}^{2}$, then it seems like the whole cell should act as one big receptor and $a_{\rm eff}\rightarrow a_{\rm cell}$. How does this work?
In a fabulous bit of hand waving plus analogies, Berg and Purcell argued that the tendency of random walk trajectories to bounce along a surface leads to correlations in the encounters of individual molecules with nearby receptors. With $N_{\rm r}$ receptors scattered over the surface of the cell, this leads to an effective size
@@
What the Berg–Purcell limit tells us is that the random arrival of molecules at their target site generates an effective concentration noise with variance
@@
Figure 24: Optimizing information about the concentration of a single transcription factor. (A) A map of $F$ from Eq (138) as a function of $K$ and $h$, showing a single clear optimum at reasonable parameters ($c_{\rm max}/c_{0}=1$). (B–F) Optimal input/output relations $\bar{g}_{\rm i}(c)$ with five target genes, with increasing values of $c_{\rm max}/c_{0}$ indicated by dashed vertical lines. At smaller $c_{\rm max}/c_{0}$ the optimal solutions are redundant, so fewer distinct input/output relations are visible. (G) Optimal values of $K/c_{0}$, showing a sequence of bifurcations as $c_{\rm max}/c_{0}$ increases and responses become distinct. Redrawn from Ref [201], with thanks to G Tkačik and AM Walczak.
@@
Figure 25: Redundancy reduction. (A) A single transcription factor (concentration $c$) activates multiple genes $\{g_{1},g_{2},g_{3}\}$ that repress one another. (B) Mean expression levels that optimize information transmission, with $c_{\rm max}/c_{0}=10$. Redrawn from Ref [202], with thanks to G Tkačik and AM Walczak.
@@
Figure 26: Schematic of direct transcriptional regulation (DTR) and indirect translational regulation (ITR). (A) In DTR, as above, activator TFs (green squares at concentration $c$) interact with binding sites along the DNA to activate expression of the regulated gene $g$. (B) In ITR scenario, input molecules (green squares) bind to mRNAs $m$ of protein $y$ (red chain) to make the mRNA inaccessible for translation (gray oval). Translation proceeds from unbound mRNA molecules, giving rise to proteins $y$ (red stars). These proteins act as repressors for gene $g$; the overall mapping from $c\rightarrow g$ is thus activating in both scenarios. Redrawn from Ref [206], with thanks to TR Sokolowski, G Tkačik, and AM Walczak.
In Figures 24B–F we see that even as the different target genes acquire distinct input/output relations, the optimal solutions still do not make much use of the lowest accessible concentrations; this regime is simply too noisy to allow reliable information transmission. But suppose that the cell makes many mRNA molecules of some intermediate protein $y$, and that the protein $c$ binds to these mRNA and represses translation, as in Fig 26; $y$ can then act indirectly as a transcriptional repressor so that the whole path $c\rightarrow g$ is activating. Then the “measurement” of low concentrations is made in parallel at many sites, rather than at just one site along the DNA, and the response is largest and most reliable at the lowest concentrations. This suggests that a molecule which functions both as a transcriptional activator and a translational repressor could make better use of its full dynamic range. It requires a careful calculation to show that information transmission really can be larger even when we constrain the total number of molecules, but it works [206]. One of the primary maternal morphogens in the fly embryo belongs to a whole class of proteins that function as such dual transcription/translation regulators [207, 208, 209, 210], and our best estimates put this system in the regime where dual regulation enhances information transmission. I found it remarkable that this almost baroque level of biological complexity can be derived as part of the solution to a fundamental physics problem faced by the cell.
Each of these pieces—repression to reduce redundancy, feedback to average over time, diffusion to average over space, and dual regulation—is an interesting problem by itself, and it was possible to make progress with a combination of analytic and (modest) numerical approaches. But all of these things are happening at once in the fly embryo, even just in the gap gene network. To put (most of) these pieces together requires a more sophisticated approach [211].
@@ 4.3 The gap genes, once more
One attractive feature of the MWC model is that it generalizes to having multiple TFs converge to regulate a single gene. Consider that gene $\rm i$ has $H_{\rm ij}$ binding sites for the protein encoded by gene $\rm j$ and $H_{{\rm i}\alpha}$ sites for the maternal input $\alpha$. Then if we follow the same equilibrium statistical mechanics arguments as above, we will find
@@
Figure 29: Maximizing positional information in the gap gene network. Random initial parameter settings produce zero positional information, with all gap genes fully on or off across the length of the embryo (1). Across ${\sim}20,000$ steps of optimization we find parameter settings that allow first one then two, three, and finally all four gap genes to be modulated by the dynamic range of maternal inputs ($2\rightarrow 3\rightarrow 4$). The majority of the optimization run then tunes the parameters to maximize the positional information, arriving at a setting ($5$) that behaves much like the real network. Redrawn from Ref [211], with thanks to T Gregor, TR Sokolowksi, and G Tkačik.
@@
  • If we eliminate feedback loops we can still reach similar levels of positional information but only with unphysically large values of the $H$ parameters; within realistic settings feedback is essential.
@@
I suspect we are just scratching the surface. Perhaps the most intriguing observation is that different local optima have broadly consistent gap gene expression patterns, but with shifts and swaps similar to those seen in related species [221, 222, 223].
@@ 5 Conclusion
As physicists we are used to the idea that macroscopic phenomena are described by coarse–grained models, and that in these models many microscopic details have disappeared (see Approach 3 in §1.2). But in the examples we understand, from the inanimate world, we can do this coarse–graining explicitly to see which (relevant) elements of the model survive to influence macroscopic dynamics and which (irrelevant) details are erased. As far as I know nobody knows how to start with a realistic description of ion channels and coarse–grain to arrive at a model of neural networks. Further, we do know that some molecular details must matter, because particular classes of cells in the brain will switch from using one kind of channel to another at crucial moments in development, and different types of cells use different combinations of channels. We also know that not all macroscopic quantities are universal, and there is the danger that what is universal might not be relevant for the organism.
@@ Acknowledgements
In addition to the handful of friends who were at the origins of these ideas, the specific things I discussed in these lectures involved collaborations with Naama Brenner, Julien Dubuis, Adrienne Fairhall, Dmitry Krotov, Geoff Lewen, Geoff Owen, Mariela Petkova, Marc Potters, Fred Rieke, Shiva Sinha, Thomas Sokolowski, and Aleksandra Walczak. This bland listing doesn’t do justice to all the fun we have had together.